OpenStudy (pleasehelpmenow):

solve the equation for a free medal!!! 4y-x=4 y+x=6

OpenStudy (anonymous):

5y = 10 y = 2 x = 4

OpenStudy (pleasehelpmenow):

please show more work

OpenStudy (anonymous):

add the two equations you get 5 y = 10 y = 2 2 + x = 6 x = 4

OpenStudy (anonymous):

another way to think about it is rearranging the second one to get x=6-y substitute this into the first equation and you get 4y-(6-y)=4 4y-6+y=4 5y=10 y=2 substitute this back into your original x equation x=6-2 x=4

OpenStudy (anonymous):

are you ok with this?